OpenStudy (anonymous):

Choose the correct simplification of the expression the (4x/y)^2. A. 4x^2/y^2 B. 8x^2/y^2 C. 16x^2/y^2 D. 16x^2/y

OpenStudy (oaktree):

Square both the top and the bottom. It should look like this:\[\frac{ (4x^2)^2 }{ y^2 } = \frac{ 16x^4 }{ y^2 }\]So answer choice C.

OpenStudy (oaktree):

Thanks for the medal! Glad to help.